Discrete Math

Use mathematical induction to prove that for all positive integers n, 2 + 4 + ... + (2n) = n(n+1).
Answer #1

let, sn= 2 + 4 + ... + (zn), nEN. Suppose. Sn = n(n+1) for some NEN [le the result is SOLA Now, for n=1 Sn=s,= 2 = 1x 2 = n(nA solution to your problem has been given above.....